Earl selects a card at random from a standard deck of cards. What is Earl's chances of selecting either a 6 or 7?
Answer by ewatrrr(24785)


Hi
Earl selects a card at random from a standard deck of cards 
(52 cards, 4 6s & 4 7s in the deck)

P(6 0r 7) = 4/52 + 4/52 = 8/52 = 2/13

Wish You the Best in your Studies.